OpenStudy (anonymous):

Given the function f(x) = 2(x + 6), find x if f(x) = 22.

OpenStudy (anonymous):

f(x) = 2(x + 6) and f(x) = 22 So, 2(x+6) = 22 2x + 12 =22 2x +12 - 12 = 22-12 2x = 10 Divide both sides by 2, what do you get?